PartnerObserving Log for 2007 Jan 22/23

Observers: Rick Pogge
Inst. Support: None
TOs: Aaron Ceranski

A winter storm dumped about a foot of fresh powdery snow on the summit last night and this morning. John Hill & Olga Kuhn are now planning to come up tomorrow given how challenging it would be to get up here today. None of the day crew has been able to make it up, compounded by snow and power outages in Safford where basecamp is shut down. We've also heard a report over the radio of about 3 feet of new snow at the end of the paved road.

The Sun came out briefly about noon so I took a brief walk down to the VATT to snap a few pictures of the winter scene. Looks like the system is moving on, so most of what we'll have to contend with is the blowing snow. It is very pretty up here.

LBT after the storm (2007 Jan 22)
LBT_20070122.jpg

Fired up LBCB at 0h UTC and ran 25bias.ob with dome lights off and the telescope zenith pointing (where it was parked last night). One of the images failed with a "too few pixels" error at UTC 2007/01/23 01:11:56.956 according to the Log Analyzer. Missing image would have been between lbc.20070123.011056 and .011203. No activity after that. Turned off LBCB

At around 0700 UTC we noticed that the dewar temperature and pressure were high (275K and 1.26mbar, respectively). See IssueTrak #442 for details. Bottom line is that the cryostat appears to have warmed up and lost high vacuum. GUI House Keeping display only plots data since about 05:00 UTC.

Linux workstation lbtmu03 logged in as sdtobs is also misbehaving in terminals (cannot do ls in any directory) See IssueTrak #443 for a description.

-- RichardPogge - 23 Jan 2007
Topic revision: r13 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback